A and B are two independent events with equal probabilities such that P(A ∪ B) = 0·75, thenA और B समान संभावनाओं वाली दो स्वतंत्र घटनाएँ हैं जैसे कि P(A ∪ B) = 0·75, तो
A और B समान संभावनाओं वाली दो स्वतंत्र घटनाएँ हैं जैसे कि P(A ∪ B) = 0·75, तो
Show answerHide answer
Correct answer: (C) P(B) = 0·5P(B) = 0·5
Explanation
Let P(A) = P(B) = p. Independence gives P(A ∪ B) = 1 − (1 − p)² = 0.75, so (1 − p)² = 0.25 and p = 0.5.मान लें P(A) = P(B) = p। स्वतंत्रता से P(A ∪ B) = 1 − (1 − p)² = 0.75, अतः (1 − p)² = 0.25 और p = 0.5।